# GRAFT: GRaPH and Table Reasoning for Textual Alignment — A Benchmark for Structured Instruction Following and Visual Reasoning

### v2

@@ 1 Introduction
Multimodal foundation models are advancing rapidly, excelling in tasks like image captioning, VQA, document parsing, and embodied reasoning. While benchmarks such as VQAv2[2], VizWiz[3], and DocVQA [4] demonstrate progress, they focus mainly on natural images or scanned documents, often lacking structural complexity or controllability. Consequently, these datasets do not fully assess models’ abilities in fine-grained, structured visual reasoning.
Although recent models perform well on general multimodal tasks, their capabilities with structured visual data—such as charts and tables—are not well understood. Tasks involving trends, relationships, or numerical reasoning in such formats require analytical skills and structured outputs (e.g., JSON, YAML), beyond basic perception.
@@ 1.1 Motivation
To address this gap, we introduce GRAFT—GRaPH and Table Reasoning for Textual Alignment—a benchmark for evaluating multimodal models on structured reasoning with visual data. GRAFT uses programmatically generated synthetic data, with all visual assets created via Python plotting libraries for precise control and reproducibility.
@@ 1.2 Research Questions
GRAFT investigates:
@@
    How well do current models handle structured visual reasoning across chart and table formats?
@@
    Can models follow complex instructions and generate correct structured outputs?
@@
    Do models differ in failure patterns between chart- and table-based inputs?
@@
    What are the limitations in visual grounding and semantic alignment?
GRAFT thus provides a controlled environment for deeper insights into model reasoning and guides future vision-language research.
The full GRAFT dataset can be found at https://huggingface.co/datasets/ServiceNow-AI/GRAFT_benchmark. We discuss additional dataset details and examples in Appendix C.
@@ 2.1 Multimodal Benchmarks
2.2 Instruction-Following in Vision-Language Models
3 Methodology
We introduce GRAFT, a fully synthetic benchmark pipeline for evaluating instruction-following and visual reasoning in multimodal large language models (MLLMs) using structured QA over charts and tables. Below, we summarize the data generation and validation process.
Figure 1: Pipeline for image generation and visual reasoning task creation.
3.1 Synthetic Visual Generation
Each GRAFT instance samples domain contexts (e.g., healthcare, finance), assigning persona and scenario metadata.
Charts: Tabular data (8–20 rows, 3–6 columns) is visualized as bar, line, pie, or scatter plots using Matplotlib/Seaborn D.1. LLM-based assessors check type fidelity and label clarity; low-quality charts are filtered out.
Tables: Tabular datasets are synthesized for various formats (pivot, grouped, sortable), then rendered as stylized images using Matplotlib D.3. LLMs assess data fidelity and labels, discarding subpar tables.
3.2 Question and Answer Generation
Approved visuals receive structured descriptions from a VLM, capturing key features. Reasoning-based questions (e.g., ranking, aggregation, anomaly detection) are then generated D.2, D.4. Each QA pair consists of:
  • A question answerable solely from the image.
  • A high-quality answer in JSON or YAML format.
Strict constraints ensure no use of external knowledge, reliance on visual labels for numeric ops, and schema-compliant formatting.
3.3 Jury-of-Judges Filtering
A Jury-of-Judges approach[31] uses multiple LLMs as independent evaluators, each scoring answers by correctness, completeness, visual grounding, and format fidelity.
Consensus strategies included majority as this filtering step ensures high-quality, reliably grounded evaluation data for benchmarking multimodal reasoning in GRAFT.
4 Experimental Setup
  • Chart-QnA: Reasoning over bar, line, scatter, and pie charts.
Table 1: Summary of GRAFT Benchmark Dataset Statistics
Chart-QnA 1,412 229 339 164 425
Table-QnA 1,739 323 551 1,022 1,620
4.1 Benchmarking Pipeline
Figure 2: Pipeline for Model Evaluation.
4.2 Evaluation Metrics
  • Correctness: Factual alignment with the visual and reference.
  • Completeness: Fully addresses the query.
  • Visual Grounding: Accurate interpretation of visual elements.
  • Format Fidelity: Strict adherence to structured output format.
5 Results
5.1 Quantitative Performance Overview
For a detailed breakdown of model performance on each subset, see Tables 8 and 9 in the Appendix.
Overall, GRAFT effectively distinguishes between shallow and deep multimodal reasoning abilities. Further discussion is provided in Section F.
6 Conclusion
7 Limitations and Future Scope
Appendix A Evaluation Criteria
To ensure consistent and transparent assessment of model responses in GRAFT, we define a scoring matrix across several key evaluation axes. Table 2 outlines the criteria for Correctness, Completeness, Visual Grounding, and Format Fidelity, with detailed interpretations for each score from 1 to 5. This framework facilitates fine-grained, aspect-based evaluation, enabling nuanced analysis of model strengths and weaknesses in structured visual reasoning tasks.
Table 2: Scoring Matrix for Evaluation Metrics (Scale: 1–5)
Evaluation Axis Score Interpretation
Correctness 5 Fully correct and factually aligned with chart/table and reference.
4 One minor factual error with negligible impact.
3 Moderate inaccuracies that weaken reliability.
2 Major factual mistake affecting the core content.
1 Entirely incorrect answer.
Completeness 5 Completely answers all parts of the question.
4 One small aspect is missing.
3 Several required components are missing.
2 Only partially answers the question.
1 Largely incomplete or missing response.
Visual Grounding 5 Perfect interpretation of visual elements.
4 Slight misread but generally correct.
3 Some visual elements misinterpreted.
2 Several major misreads of visual content.
1 No clear connection to the visual data.
Format Fidelity 5 Output matches schema and formatting exactly.
4 Minor format issues (e.g., indentation, punctuation).
3 Multiple format errors, still readable.
2 Major format breakdown (e.g., broken structure).
1 Completely incorrect or unusable format.
Appendix B Benchmark Comparison
To contextualize GRAFT within the broader landscape of multimodal evaluation, Table 3 provides a comparative overview of widely used benchmarks and highlights the unique innovations introduced by GRAFT. For each benchmark, we summarize its core focus and describe how GRAFT advances the field through new capabilities, greater control, or expanded evaluation dimensions. This comparison underscores GRAFT’s role in addressing the limitations of existing datasets and setting a higher standard for structured visual reasoning and output alignment.
@@ Appendix B Benchmark Comparison
Appendix C Dataset Preview
The following appendix provides a visual walkthrough of sample records drawn from the GRAFT benchmark. Each record corresponds to a visual reasoning QA instance—grounded in either a chart or a table—and includes both the visual input and detailed annotations used during training and evaluation.
Each data record is composed of the following key fields:
  • id: A unique SHA-256 hash representing the data instance.
  • image: The rendered chart or table image used as the sole visual input to the model.
  • conversation: A structured list capturing the instruction-response exchange. This includes the visual reference (image) followed by a complex, multimodal question and a YAML-structured answer generated by a VLM.
  • metadata: Encapsulates the quality and validity judgments for each record, including:
    • answerability.JUDGEMENT: Whether the question is visually answerable.
    • answerability.JUDGEMENT _EXPLANATION: Explanation for the answerability tag.
    • correct_answer_evaluation.JUDGEMENT: Whether the answer is factually and visually correct.
    • chosen_formattedness.JUDGEMENT: Assessment of whether the answer adheres to the expected YAML/JSON format.
    • generated_instruction, generated_answer: The actual model-generated instruction and structured output.
    • image_description: Auto-generated natural language description of the visual, capturing axes, legends, trends, and key patterns.
  • judgement_metadata: A consolidated set of scores and explanations from multiple model-based judges (e.g., GPT-4o, Pixtral, Mistral), including their independent verdicts and rationales. This field helps support ensemble-based QA validation pipelines described in Section 4.1.
Each previewed record in the pages below is presented as a compact tabular summary with its associated chart or table image. Fields with long content are truncated for readability (indicated by ellipses).
Sample rendered records are shown below for the first five examples in the benchmark:
Appendix D Instruction Categories
This appendix summarizes the instruction categories used in MIRAGE for generating diverse question types across different visual modalities: graphs, tables, natural images, and posters.
D.1 Graph Generation Taxonomy
Graph-based questions in GRAFT are grounded in high-quality, programmatically rendered charts created from synthetic domain-specific data. The generation process ensures diverse visual styles, realistic semantics, and domain-grounded personas. TableD.3 outlines the key dimensions used for chart synthesis.
Table 4: Taxonomy of Graph Image Generation Attributes
Dimension Description and Examples
Domain Genre Drawn from 25+ application areas including Finance, Education, Environment, and Cybersecurity. Each genre is associated with realistic user roles such as researchers, planners, or policy makers.
Geographic Context Sampled from over 50 global cities and countries to provide cultural and locational grounding. Examples include: Nairobi, Tokyo, San Francisco, and Dubai.
Plot Type Determines visual encoding style. Includes: line, bar, scatter, hist, kde box, heatmap, pie, area, violin count, strip, swarm Each type is selected based on the semantics of the underlying data (e.g., trends, distributions, comparisons).
Visual Styling Rendered using Python plotting libraries (e.g., Matplotlib, Seaborn) with variation in: Axis labels, tick density, color palettes Title fonts and padding DPI scaling (150–350) for clarity and size diversity
Semantic Context Each graph is grounded in a structured metadata schema, which includes: Persona (e.g., high school teacher, financial analyst) Scenario (e.g., monthly earnings report, student performance review) Timeframe and geographic reference
This comprehensive taxonomy ensures the GRAFT graph dataset supports wide-ranging reasoning tasks grounded in diverse, plausible, and richly contextualized scenarios.
D.2 Graph Instruction Categories
Graph-based QA tasks are constructed to probe core visual reasoning capabilities over structured chart imagery. Each question type targets a distinct reasoning skill, grounded solely in the information visible within the chart.
Table 5: Instruction Taxonomy for Graph-Based QA
Category Description Example Question
Comparison Compare values between groups or categories Which group had the highest value?
Trend Identification Identify patterns, slopes, or growth/decline trends What trend is observed from 2010 to 2020?
Change Calculation Determine difference over time or between points By how much did revenue grow from Q1 to Q2?
Ranking Order categories based on visual magnitude Rank regions by number of cases reported
Proportion/Share Determine parts of a whole or segment sizes What share of traffic is from mobile users?
Outlier Detection Identify anomalies or deviations Which year shows an unusual spike in complaints?
Inference Multi-step reasoning from visual data If the current trend continues, when will value exceed 1,000?
Aggregation Sum or combine values from a plot What is the combined total of Category A and B?
Ratio/Rate Compute visual ratios or per-unit rates What is the dropout rate per 100 students?
This instruction taxonomy ensures systematic coverage of reasoning capabilities expected in data visualization tasks and enables benchmarking of visual-language models on structured visual inputs.
D.3 Table Generation Taxonomy
To support structured visual reasoning in the GRAFT benchmark, we design a comprehensive taxonomy for synthetic table generation. This governs not only the content and semantics of tables but also their visual formatting and contextual grounding. Table 6 summarizes the core dimensions of variation used to produce high-fidelity table images.
Table 6: Taxonomy of Table Image Generation Attributes
Dimension Description and Examples
Domain Genre Sampled from 25+ domains (e.g., Healthcare, Finance, Technology) each tied to realistic personas like analysts, doctors, teachers, or planners.
Geographic Context Includes real-world city/country names (e.g., Tokyo, Cape Town, London) to simulate plausible contexts.
Table Type Encodes layout structure. Categories include: flat, grouped, pivot, sortable, highlighted comparison, ranked, time_series, proportion, matrix
Visual Styling Controlled variation in: Font families (e.g., Arial, STIXGeneral), font sizes (8–14pt) Header and row colors, edge borders (black/gray variants) Column/row scaling, zebra striping, padding Render DPI (150–350) for visual fidelity
Layout Controls Additional variation in: Minimum width/height constraints Title font and spacing Highlight rules for emphasizing key cells
This generation taxonomy ensures that each table is semantically grounded, visually distinct, and structurally diverse—reflecting the range of layouts seen in real-world analytical tools and dashboards.
D.4 Table Instruction Categories
Table-based QA tasks are constructed to probe various reasoning capabilities over structured visual data. We define a core taxonomy of instruction categories to ensure analytical diversity across the benchmark. Table7 summarizes these categories with descriptions and example question types.
Table 7: Instruction Taxonomy for Table-Based QA
Category Description Example Question
Comparison Compare values across rows or columns Which group has the highest value?
Pattern Recognition Detect structural or numeric patterns Which column shows a consistent increase?
Change Calculation Compute differences over time or category What changed most between Year X and Y?
Ranking Sort entries based on metrics Rank departments by average score.
Proportion/Share Calculate part-to-whole relationships What % of revenue comes from Product A?
Anomaly Detection Identify unusual highs/lows Which month saw an unexpected drop?
Deductive Inference Multi-step logic or scenario-based reasoning If salaries rise by 10%, what’s the total cost?
Aggregation Sum or average values across a dimension What is the total enrollment in Q3?
Ratio/Rate Derive ratios, per-capita, or unit rates What is the conversion rate per 100 visits?
These QA categories ensure comprehensive coverage of the reasoning spectrum required for structured data interpretation, enabling targeted evaluation of visual understanding and analytical reasoning.
Appendix E Detailed Model Results
E.1 Table-QnA Results
Table 8: Evaluation scores on the Table-QnA subset of GRAFT (1,739 examples).
Model Correctness Completeness Visual Grounding Format Fidelity Avg. Score
Qwen 2.5 32B VL 4.20 4.90 3.85 4.98 4.48
GPT4o 3.78 4.91 3.46 4.97 4.28
Mistral 24B 3.76 4.91 3.44 4.98 4.27
Pixtral 12B 3.40 4.88 3.17 4.99 4.11
GPT4o-mini 3.39 4.80 3.20 4.97 4.09
E.2 Chart-QnA Results
Table 9: Evaluation scores on the Chart-QnA subset of GRAFT (1,412 examples).
Model Correctness Completeness Visual Grounding Format Fidelity Avg. Score
Pixtral 12B 3.92 4.90 3.37 4.99 4.29
Qwen 2.5 32B VL 3.92 4.90 3.34 4.99 4.29
Mistral 24B 3.86 4.94 3.25 4.99 4.26
GPT4o 3.83 4.88 3.25 4.99 4.24
GPT4o-mini 3.62 4.87 3.08 4.99 4.14
Appendix F Model Results Discussion
This section presents a deeper qualitative analysis of model behavior on the GRAFT benchmark, considering architectural, training, and alignment differences. We also integrate insights from our evaluation to highlight where models succeed or fail, and how this reflects their design choices. A synthesis of these findings follows in the form of general takeaways.
F.1 Model-Specific Observations
Qwen-2.5 32B VL demonstrates consistent superiority across all major dimensions—correctness, grounding, and structure. Its high correctness scores across both Chart-QnA and Table-QnA subsets suggest strong internal alignment with step-by-step reasoning, likely supported by extensive instruction tuning and multimodal pretraining. The model’s stable behavior across modalities also indicates robustness in both spatial (charts) and relational (tables) comprehension.
Pixtral 12B is a standout performer on Chart-QnA, attaining the highest visual grounding and average score. However, its performance drops sharply on Table-QnA, particularly in correctness. This dichotomy suggests Pixtral’s strengths lie in shallow visual-text alignment and layout parsing—useful in charts—but it lacks the depth required for multi-step symbolic reasoning over tabular data. This is consistent with its training objectives, which prioritize efficiency and image-level instruction following over structured reasoning.
Mistral-Small-24B-Instruct performs uniformly well across both modalities, especially in format fidelity and completeness. It shows reliable but slightly underwhelming correctness on tables and charts compared to Qwen. Mistral’s performance likely reflects its balanced instruction tuning and decoding stability. However, the absence of a visual modality natively built into Mistral may explain its modest visual grounding scores, suggesting dependence on weak cross-modal adaptation layers.
GPT4o displays strong formatting and completeness scores, which align with OpenAI’s focus on controllability and schema compliance. However, the model consistently underperforms on visual grounding and correctness—especially on complex Chart-QnA questions. This supports the interpretation that GPT4o’s visual capabilities, while promising, are still grounded more in OCR and text layout heuristics than deep visual-semantic parsing. The disparity between visual and textual modalities may account for misalignments in compositional reasoning.
GPT4o-mini, a smaller and more efficient variant, ranks lowest on most axes, particularly correctness and grounding. Despite strong formatting, the model struggles with semantic interpretation of structured data. Its performance illustrates the trade-off between lightweight deployment and high reasoning fidelity. Given its compressed size and likely limited vision training, this result is expected and confirms the importance of scale in visual instruction tuning.
F.2 Analysis and Key Takeaways
These results underscore several crucial observations:
  • Correctness is non-negotiable. No matter how complete or well-formatted a response is, an incorrect answer undermines the purpose of the task. Models with higher correctness scores should be prioritized over those with marginally higher averages but weak reasoning capabilities.
  • Format fidelity is uniformly strong across all models, highlighting the success of schema-constrained instruction templates. This suggests that even smaller models can be guided to produce well-structured outputs when appropriately prompted.
  • Visual grounding remains a bottleneck, particularly for table-based questions that involve implicit relationships, long-range dependencies, or hierarchical reasoning.
  • Instruction tuning and scale matter. Larger instruction-tuned models like Qwen and Mistral outperform others across correctness and grounding, reflecting better task alignment and internal representation.
These insights are further illustrated in the accompanying visual summaries, providing a holistic view of model strengths and limitations across modalities.

### v3

@@ Abstract
Keywords large language models $\cdot$ multimodal benchmarks $\cdot$ instruction following $\cdot$ visual reasoning $\cdot$ chart understanding $\cdot$ table reasoning $\cdot$ visual-textual alignment $\cdot$ synthetic data generation $\cdot$ evaluation framework
@@ 1 Introduction
Recent years have witnessed rapid advancements in multimodal foundation models, which have achieved impressive performance across diverse tasks such as image captioning, visual question answering (VQA), document understanding, and embodied reasoning. Widely used benchmarks, including VQAv2 [2], VizWiz [3], and DocVQA [5], have played a critical role in quantifying progress in these domains. However, these datasets are predominantly oriented toward natural images or scanned documents and often lack either structural complexity or controllable variability. As a result, they provide limited insight into models’ abilities to perform fine-grained reasoning over structured visual content.
Although contemporary multimodal systems excel at broad perceptual and reasoning tasks, their performance on structured visual data—particularly charts and tables—remains insufficiently explored. Reasoning over such inputs requires not only perceptual recognition but also higher-order analytical competencies, such as identifying trends, comparing relationships, and performing numerical inference. Moreover, many practical applications demand structured outputs (e.g., JSON, YAML) that extend beyond natural language responses, thereby introducing further challenges in alignment between visual understanding and structured generation.
@@ 1.1 Motivation
To bridge this gap, we present GRAFTGRaPH and Table Reasoning for Textual Alignment—a benchmark explicitly designed to evaluate multimodal models on structured reasoning tasks involving charts and tables. GRAFT leverages programmatically generated synthetic datasets, with all visual assets produced using Python plotting libraries to ensure reproducibility, fine-grained control, and systematic variation in task complexity. This design enables a more rigorous evaluation of reasoning capabilities, disentangled from the noise and biases inherent in natural datasets.
@@ 1.2 Research Questions
GRAFT is developed to investigate the following key research questions:
@@
    To what extent can current multimodal models perform structured reasoning over chart- and table-based inputs?
@@
    Are models capable of following complex, multi-step instructions and generating semantically and syntactically correct structured outputs?
@@
    Do error and failure patterns differ systematically between chart reasoning and table reasoning tasks?
@@
    What are the existing limitations of multimodal models in terms of visual grounding and semantic alignment with textual instructions?
By addressing these questions, GRAFT provides a controlled and transparent environment for diagnosing the reasoning abilities of multimodal systems. The benchmark thereby offers deeper insights into existing limitations and establishes a foundation for advancing vision-language research toward robust structured reasoning.
The full GRAFT dataset is publicly available on Hugging Face [^1] https://huggingface.co/datasets/ServiceNow-AI/GRAFT_benchmark. Additional details regarding dataset construction, along with illustrative examples, are provided in Appendix Dataset Preview.
@@ 2.1 Multimodal Benchmarks
To contextualize GRAFT within the broader landscape of multimodal evaluation, Table 1 provides a comparative overview of widely used benchmarks and highlights the unique innovations introduced by GRAFT.
@@
2.2 Instruction-Following in Vision-Language Models
3 Methodology
We introduce GRAFT, a fully synthetic benchmark pipeline for assessing instruction-following and structured visual reasoning in multimodal large language models (MLLMs) over charts and tables. The pipeline comprises three stages: (i) synthetic visual generation, (ii) structured QA construction, and (iii) multi-model quality control via a jury-of-judges filter. Figure 1 provides a high-level overview. A complete reference of the variables and notation introduced in this section is provided in Appendix Notation and Variable References.
Figure 1: GRAFT pipeline for synthetic visual generation and structured reasoning task construction.
3.1 Problem Setup and Notation
Let $v$ denote a visual instance (chart or table image), $d(v)$ a structured description of $v$, $q$ a reasoning query grounded in $v$, and $a$ a structured answer adhering to a schema $\mathcal{S}$ (JSON or YAML). An MLLM $f_{\theta}$ receives $(v,q)$ and must produce $\hat{a}\in\mathcal{S}$ such that $\hat{a}$ is semantically correct and visually grounded.
Input Spaces.
Domains $\mathcal{D}$ (e.g., healthcare, finance), personas $\mathcal{P}$, and scenarios $\mathcal{X}$ control content and style; chart types $\mathcal{C}$ and table types $\mathcal{T}$ control structure (Appendices Graph Generation Taxonomy, Table Generation Taxonomy).
Output Targets.
For each instance we produce $(v,d(v),q,a,\mathsf{meta})$, where $\mathsf{meta}$ stores lineage (random seeds, generators, versioned prompts), and validators’ scores.
3.2 Stage I: Synthetic Visual Generation
Inputs.
Domain $d\in\mathcal{D}$, persona $p\in\mathcal{P}$, scenario $x\in\mathcal{X}$; tabular seeds $T\in\mathbb{R}^{m\times n}$ (with $m\in[8,20],\,n\in[3,6]$); chart/table configuration $\kappa$.
Outputs.
A validated visual $v$ (PNG), its rendering specification $\rho$ (Matplotlib/Seaborn parameters), and metadata $\mathsf{meta}_{\text{viz}}$ (axes labels, legends, color maps, font sizes, DPI, seeds).
Procedure.
  1. 1.
    Scenario sampling. Sample $(d,p,x)$ to bind narrative context and label spaces (units, categories).
  2. 2.
    Data synthesis. Generate tabular seed $T$ with typed columns (categorical, ordinal, numerical), injected with controllable traits (trend magnitude, noise level, sparsity) via parameterized generators.
  3. 3.
    Chart rendering. For charts, select $c\in\mathcal{C}$ (bar/line/pie/scatter). Render $v$ via Matplotlib [^2] https://matplotlib.org//Seaborn [^3] https://seaborn.pydata.org/ using $\rho$; enforce axis titles, tick formatting, legends, and numeric precision (Appendix Graph Generation Taxonomy).
  4. 4.
    Table rendering. For tables, select $t\in\mathcal{T}$ (pivot, grouped, sortable); format headers, row bands, and cell alignment; render to image with Matplotlib (Appendix Table Generation Taxonomy).
  5. 5.
    LLM-based visual validation. Use automatic assessors to rate: (i) type fidelity (plot/table matches schema), (ii) label clarity (axes/headers/legend legible, units present), (iii) readability (no occlusion/overplotting). Reject on any failing criterion; otherwise persist $(v,\rho,\mathsf{meta}_{\text{viz}})$.
3.3 Stage II: Structured QA Construction
Inputs.
Validated visual $v$ with spec $\rho$; domain context $(d,p,x)$; instruction-type taxonomies $\Gamma_{\text{chart}}$ and $\Gamma_{\text{table}}$ (e.g., ranking, aggregation, comparison, anomaly detection; Appendices Graph Instruction Categories, Table Instruction Categories); output schema $\mathcal{S}$.
Outputs.
A tuple $(d(v),q,a)$ where $d(v)$ is a structured description, $q$ is a visual-grounded question answerable solely from $v$, and $a\in\mathcal{S}$ is a machine-parseable answer.
Constraints.
(i) No external knowledge: all facts must be derivable from labels or marks in $v$; (ii) Numeric grounding: computations must reference chart/table labels explicitly; (iii) Schema validity: $a$ must satisfy $\mathcal{S}$ (field names, types, enumerations).
Procedure.
  1. 1.
    Saliency extraction. A VLM summarizes $v$ into $d(v)$ capturing encodings (marks, channels), key trends (slopes, extrema), category relations, and plausible aggregations.
  2. 2.
    Instruction sampling. Sample a reasoning template $\gamma\in\Gamma_{\text{chart}}\cup\Gamma_{\text{table}}$ conditioned on $(d,p,x)$ and the structure in $d(v)$.
  3. 3.
    Question synthesis. Instantiate $q$ from $\gamma$ ensuring answerability from $v$ alone; inject clarifying spans (units, time ranges, category subsets) to prevent ambiguity.
  4. 4.
    Answer derivation. Compute $a$ by programmatic reading of $T$ (the generator’s source of truth) and by an independent VLM read from pixels; require agreement. Serialize $a$ to $\mathcal{S}$ (JSON/YAML) with strict formatting.
  5. 5.
    Schema and determinism checks. Validate $a$ against $\mathcal{S}$; re-run derivation with fixed seeds to ensure deterministic reproduction; reject if mismatched.
3.4 Stage III: Jury-of-Judges Filtering
Inputs.
Candidate triples $(v,q,a)$ with $d(v)$ and $\mathsf{meta}$.
Outputs.
A curated set $\mathcal{G}$ of accepted instances; per-instance diagnostics (scores, rationales, failure tags).
Scoring Dimensions.
(i) Correctness (semantic equivalence to ground truth), (ii) Completeness (all requested fields present), (iii) Visual grounding (answer aligns to marks/labels in $v$), (iv) Format fidelity (valid $\mathcal{S}$) [32].
Procedure.
  1. 1.
    Independent evaluation. $K$ LLM judges score each $(v,q,a)$ along the four dimensions with calibrated rubrics (Likert or binary per field) and short textual justifications.
  2. 2.
    Consensus aggregation. Apply majority vote for accept/reject and average for continuous rubrics; tie-break with a deterministic priority (visual grounding $\rightarrow$ correctness $\rightarrow$ format).
  3. 3.
    Adjudication & repair. For borderline cases, attempt auto-repair (format-only or minor rounding) under constraints that forbid semantics changes; re-validate schema; otherwise reject.
  4. 4.
    Finalization. Persist accepted instances to $\mathcal{G}$ with full lineage: $(v,\rho,T,d,p,x,d(v),q,a,\text{scores},\text{judge\_notes},\text{seeds})$.
The GRAFT pipeline combines programmatic data generation, structured reasoning tasks, and rigorous quality control into a unified framework for benchmarking multimodal reasoning. By leveraging synthetic visual assets, schema-constrained QA generation, and multi-model adjudication, GRAFT produces a reproducible and reliable benchmark. Together with the variable reference table in Appendix Notation and Variable References, this methodology provides a transparent foundation for evaluating and diagnosing the structured reasoning capabilities of multimodal models.
4 Experimental Setup
For all experimentation and dataset construction presented in this paper, we employed the SyGra [^4] https://github.com/ServiceNow/SyGra framework [4] as the primary tool for data generation. SyGra’s modular, graph-based architecture enables scalable synthesis, quality tagging, and management of multimodal synthetic data, powering the diverse visual–text datasets used throughout our study.
  • Chart-QnA: Reasoning over bar, line, scatter, and pie charts.
Table 2: Summary of GRAFT Benchmark Dataset Statistics
Chart-QnA 1,412 229 339 164 425
Table-QnA 1,739 323 551 1,022 1,620
4.1 Benchmarking Pipeline
Figure 2: Pipeline for Model Evaluation.
4.2 Evaluation Metrics
  • Correctness: Factual alignment with the visual and reference.
  • Completeness: Fully addresses the query.
  • Visual Grounding: Accurate interpretation of visual elements.
  • Format Fidelity: Strict adherence to structured output format.
5 Discussion
We conduct a qualitative analysis of model performance on GRAFT, focusing on architectural, training, and alignment factors. The goal is to identify where models succeed or fail, and how these outcomes reflect their underlying design choices. We conclude with general takeaways distilled from these observations.
5.1 Model-Specific Observations
Qwen-2.5 32B VL[38] achieves the most consistent performance across correctness, grounding, and structure. Its strong results on both Chart-QnA and Table-QnA indicate robust reasoning alignment, likely driven by extensive multimodal pretraining and instruction tuning. The stability across modalities suggests well-integrated spatial (chart) and relational (table) representations.
Pixtral 12B[39] excels on Chart-QnA, obtaining the highest visual grounding scores, but its accuracy degrades on Table-QnA. This divergence suggests that Pixtral’s strengths lie in shallow visual-text alignment and layout parsing—sufficient for charts—but it lacks the symbolic reasoning depth required for tables. The pattern aligns with training objectives optimized for image-level efficiency rather than structured multi-step reasoning.
Mistral-Small-24B-Instruct [^5] https://huggingface.co/mistralai/Mistral-Small-24B-Instruct-2501 delivers balanced performance across modalities, with particular strength in format fidelity and completeness. Its correctness scores, while solid, lag behind Qwen. The absence of a native visual backbone may explain weaker grounding, with performance likely mediated by cross-modal adaptation layers rather than deeply integrated vision modules.
GPT4o[37] demonstrates high format compliance and completeness, reflecting strong alignment with schema-driven prompting. However, it underperforms on correctness and visual grounding, especially for Chart-QnA. This suggests reliance on OCR-style heuristics rather than genuine visual-semantic parsing, which limits compositional reasoning across visual and textual modalities.
GPT4o-mini, the smaller variant, ranks lowest overall, particularly on correctness and grounding. While it maintains schema fidelity, it struggles with semantic interpretation of structured data. These results illustrate the trade-off between deployment efficiency and reasoning capacity, reinforcing the role of scale in effective multimodal alignment.
5.2 Key Takeaways
Our analysis highlights four central findings:
  • Correctness is paramount. Format fidelity and completeness are insufficient if the underlying reasoning is flawed. Models that consistently produce correct answers should be prioritized over those with superficially higher aggregate scores.
  • Schema adherence is largely solved. All models reliably produce structured outputs under schema constraints, indicating that prompt-based guidance is effective even for smaller models.
  • Visual grounding remains the main bottleneck. Table-based reasoning in particular exposes weaknesses in capturing implicit relations, hierarchical structures, and long-range dependencies.
  • Instruction tuning and scale drive performance. Larger, instruction-tuned models (e.g., Qwen, Mistral) achieve better correctness and grounding, underscoring the importance of both capacity and targeted alignment in structured reasoning tasks.
Together, these findings outline the current frontier of multimodal reasoning: while structural compliance is achievable, correctness and grounding—especially over relational data—remain unresolved challenges.
6 Results
6.1 Quantitative Performance Overview
For a detailed breakdown of model performance on each subset, see Tables 9 and 10 in the Appendix.
Overall, GRAFT effectively distinguishes between shallow and deep multimodal reasoning abilities. Further discussion is provided in Section 5.
7 Conclusion
8 Limitations and Future Scope